If z + 1/z=2 cos theta , theta belongs to R.  Show that |z|=1 and for any n belongs to Z, z^n + 1/z^n  =2 cos n theta.


1
Expert's answer
2021-06-18T12:30:38-0400

We note "|z|=r". With this notation the imaginary part of "z+1\/z" is "(r-1/r) \sin\theta". Now as "2\cos \theta" is a real number, its imaginary part is zero. Therefore there are two cases :

  1. "r-1/r=0 \rightarrow r=|z|=1"
  2. "\sin \theta = 0 \rightarrow z\in \mathbb{R}" but in this case we have "z=\pm 1, |z|=1"

Therefore, in any case we have "|z|=1" and so "z" can be written as "e^{i\theta}". Using this expression we find that for any "n\in \mathbb{Z}" we have

"z^n + 1/z^n = e^{in\theta} + e^{-in\theta}=2\cos n\theta"